### v1

@@ Abstract
Solar energy is one of the most promising renewable energy resources. Forecasting photovoltaic power generation is an important way to increase photovoltaic penetration. However, the task of photovoltaic forecasting is complicated due to its property of uncertainty, especially in specific regions during the foggy winter. This paper proposes a novel model to accomplish the problem. A developed entropy is created to qualify the uncertainty during the foggy winter. The clustering method and modified retention network are applied to reduce complexity and forecast, respectively. We adopt an optimization to optimize the hyperparameters. Results are validated from the multivariate forecasting model using the dataset from a photovoltaic power station in Jiangsu Province, China. Experiments show that the proposed model improves the forecasting accuracy compared to various models during the foggy winter.
@@ 1 Introduction
Fossil fuels extracted over the past few decades to meet global energy demands, causing serious environmental problems [1]. In contrast, solar power is recognized globally for its clean, pollution-free, and sustainable benefits [2]. Fig 1 shows a significant increase in solar energy capacity over the past decade, with data derived from the International Renewable Energy Agency [3].
@@
Photovoltaic (PV) power generation embraces the property of uncertainty due to external factors [4]. Consequently, this uncertainty makes forecasting accuracy decrease. The power system may lead to serious failures if inaccurate PV forecasting is applied [5]. Consequently, accurate forecasting is essential to improve the stability of the power system [6].
@@
Existing PV forecasting models can be classified into physical, statistical, and hybrid models [7]. Physical models focus on theoretically explaining the works of PV power systems, while statistical models forecast output through data analysis. Hybrid models combine multiple models to improve accuracy. Table 1 lists information on existing PV forecasting models, highlighting their descriptions, advantages, and disadvantages.
The frequent occurrence of fog poses a significant challenge to the forecasting accuracy of PV power generation. Fog particles scatter sunlight, resulting in a reduction amount of solar radiation reaching the ground [8]. This phenomenon is particularly pronounced in winter because the conditions of low temperatures and humidity are more conducive to the formation and persistence of fog. Hence, the uncertainty is amplified in foggy winter conditions, making the forecasting task more complicated. It is essential to qualify the uncertainty and make precise forecasts during the foggy winter.
@@
Existing works rarely consider the impact of fog on PV forecasting. Therefore, we create developed entropy to qualify the uncertainty affected by fog. Entropy is a way of qualifying the uncertainty in time series [9]. Then, we use the clustering method based on developed entropy to reduce complexity. Lastly, we use modified retention network (RetNet) to forecast. Drawing from the analysis above, the main contributions of this work are presented as follows:
@@
  • We create the Tsallis Entropy by Weighted Permutation Pattern (TEWPP) to qualify the uncertainty during the foggy winter. TEWPP qualifies the uncertainty by examining the distribution of patterns, reflecting the distribution for a series of PV power generation during the foggy winter.
@@
  • We use hierarchical clustering based on entropy to reduce the complexity of the model. This strategy can improve the similarity of elements in the same cluster, thereby reducing the complexity.
@@
  • We use CNN-RetNet to extract the features and forecast. CNN uses kernel on input data to extract local features through multiplication and summation. RetNet is used to forecast. The model improves the forecasting accuracy.
@@
  • We adopt the non-dominated sorting genetic algorithm-II (NSGA-II) to optimize the hyperparameters. NSGA-II searches for optimal solutions by simulating natural selection and genetic mechanisms. The robustness of the model can be enhanced by using this method.
@@
  • We use the data from a PV power station in Jiangsu, China, during winter. The experimental results are evaluated through evaluation metrics, and compared these results to others obtained from traditional models during the foggy winter.
@@
The remainder of this paper is organized as follows. Section 2 presents an overview of related works. Section 3 introduces the optimal goal of the system model. Section 4 gives an architecture and a detailed description of CCRetNet. The performance of CCRetNet is evaluated through experiments and analysis, as detailed in section 5. Section 6 provides conclusions.
@@
Physical models Simulate the physical process of PV power plants to forecast. Provide high-accuracy forecasts. Require large computational cost for measuring and updating data.
@@
Hybrid models Combine multiple models according to the characteristics of forecasting models to improve accuracy and reliability. Overcome limitations of a single model, improving accuracy and reliability. Result in higher computational cost than a single model.
@@ 2 Related works
Weather factors in the field of PV power generation have been taken into account [13], [14]. Wang et al. [15] introduced weather classification to predict PV power generation. The model could dynamically adapt to weather variations, improving the accuracy of prediction. However, the performance degraded when the weather changed significantly. Michael et al. [16] proposed a model using an optimized deep long short term memory neural network. The model demonstrated optimal performance under various meteorological conditions with univariate and multivariate data. The complexity of the model required substantial computational resources. Sun et al. [17] presented a probabilistic solar power forecasting framework based on weather scenarios. The weather scenarios through gaussian mixture model, copula, and gibbs sampling. These scenarios were input into a machine learning model to forecast. Bellinguer et al.[18] integrated weather information into the prediction model. Weather information was used as a different ensemble strategy for explanatory and state features, studying its impact on linear and nonlinear models. The proposed model significantly improved the accuracy of predictions However, the model structure became more complex due to integrating multiple information sources.
Nowadays, the attention mechanism is used in the field of PV forecasting [19]. Yu et al. [2] used improved complementary empirical mode decomposition and variational mode decomposition for reducing noise. The attention mechanism focused on critical information to improve accuracy. The model was adjusted to data features to improve accuracy, whereas the complexity of the model limited its application. Yu et al. [10] employed temporal, frequency, and fourier attention, capturing detailed time-series insights from PV power data. The attention mechanism focused on crucial temporal and environmental features, thereby enhancing accuracy. However, this approach also obscured the significance of the time series itself. Wang et al. [20] combined a temporal convolutional network with an improved deep residual shrinkage network. The special attention mechanism in the model enhanced the ability to extract features in noisy environments. However, the model required significant computational resources.
Table 2 summarizes the main related works, including details of works, advantages, and disadvantages, where "Ref." denotes "References".
@@
PV forecasting considering weather conditions
@@
[21] 2024 Weather clustering; Dual attention mechanism; Transfer learning; Target domain adaptation; Structure transferring Improved the data quality of subsequent model training. Relied on data distribution analysis for interval setting.
[22] 2024 Long-term time series forecasting; Weather forecast data; Long-short term memory; Self-attention; Time series embeddings Utilized weather forecasting data to improve forecasting accuracy effectively. Ignored how weather forecasting errors affect PV forecasting.
[23] 2023 Weather-type classification; Weather-type forecast; Symbol-sequence histograms clustering; Phase-space reconstruction; Extremely learning machine; Photovoltaic output prediction Removed noise from historical PV output data, enhancing the accuracy. Ignored the changing weather conditions throughout the whole day.
[24] 2023 Solar photovoltaic power forecast; Artificial neural network; Numerical weather prediction; Weather observational data; Physical model Processed different types of input data. Introduced a risk of data drift.
[25] 2023 Probabilistic solar forecasting; Ensemble NWP; Physical model chains; Photovoltaic power; Quantile regression; European Centre for Medium-Range Weather Forecasts Indicated the robustness and didn’t rely on a specific optimization, making it versatile for various user needs. Limited by the quality of numerical weather prediction data.
[26] 2022 Solar irradiation; Confidence interval forecast; Intra-hour horizon; Smart control; Photovoltaic generation output power Performed well under partly cloudy, cloudy, and clear weather conditions. Required to fit model parameters through a huge database.
@@
[27] 2024 Regional PV power forecast; Improved time-series dense encoder; Graph attention network Improved regional PV power prediction using multi-source spatio-temporal data fusion. Degraded model performance due to insufficient historical data.
[28] 2023 Photovoltaic power forecast; Channel attention mechanism; Lightweight networks; Time series forecasting Improved accuracy. Available for multivariate time series forecasting tasks. Caused a risk of overfitting.
[29] 2023 Photovoltaic; Dual stream network; CNN; GRU; Solar power forecasting; Renewable energy; Self-attention mechanism; CNN-LSTM; CNN-GRU Learned two spatio-temporal features simultaneously and extracted them by processing historical data in parallel. Required large computational cost.
[30] 2023 Solar photovoltaic power; Short-term prediction; Deep learning; Attention net; Bayesian optimization Enhanced processing capabilities through attention mechanisms. Sensitive to quality of input data.
[31] 2023 Short-term solar power generation prediction; Decomposition and ensemble; Sequence to sequence model; Attention mechanism; Bayesian optimization Validated perform well on a real-world dataset. Contained a lot of layers, resulting in a large amount of calculation.
[32] 2023 Long and short-term memory; Long and short term time series network; Temporal convolutional attention neural networks; PV power Forecasting Captured the spatio-temporal features of PV power and improved the accuracy. Only represented the ultra-short-term prediction.
[33] 2022 Deep attention; ConvLSTM network; Deep learning; Photovoltaic power forecast; Solar energy; Kernel density estimation; Renewable energy Improved forecasting accuracy in each season of the year. Relied on quality of data, affecting accuracy.
@@
ARIMA Autoregressive Integrated Moving Average model MSR Multi-Scale Retention
CCRetNet Clustering and CNN-RetNet NSGA-II Non-dominated Sorting Genetic Algorithm-II
CNN Convolutional Neural Network PV Photovoltaic
FLOPs Floating Point Operations Ref. Reference
FFN Feed-Forward Network RetNet Retention Network
GRU Gate Recurrent Unit RMSE Root Mean Squared Error
LSTM Long Short Term Memory RS Random Search
MAE Mean Absolute Error TPE Tree-structured Parzen Estimator
MLP Multilayer Perceptron TEWPP The Tsallis Entropy by Weighted Permutation Pattern
@@ 3 Problem definition
Assume $k$ denotes the optimal number of clusters of PV power generation. $\mathcal{L}_{n}(\cdot)$ denotes the loss function of the forecasting model for the $n$-th cluster, where $\phi_{n}$ denotes the set of parameters and $\psi_{n}$ denotes the set of hyperparameters for the $n$-th model. $\Phi=\{\phi_{1},\phi_{2},\cdots,\phi_{k}\}$ denotes the set of parameters for all model and $\Psi=\{\psi_{1},\psi_{2},\cdots,\psi_{k}\}$ denotes the set of hyperparameters for all model.
During the training phase, the objective function aims to minimize the loss function for each cluster. The objective function denotes as follows:
@@
This phase finds the optimal set $\phi_{n}^{*}$ when equation (1) reaches the minimum value (i.e. $\phi_{n}^{*}=\arg\min_{\Phi}\sum_{n=1}^{k}\mathcal{L}_{n}(\phi_{n})$).
For the validating phase, it should be assigned to the $n$-th cluster that minimizes the loss function for the new data segment denoted by $\boldsymbol{x}$. The objective function denotes as follows:
@@
This phase finds the optimal set $\psi_{n}^{*}$ when equation (2) reaches the minimum value (i.e. $\psi_{n}^{*}=\arg\min_{\Psi}\sum_{n=1}^{k}\mathcal{L}_{n}(\psi_{n})$).
@@
where $\alpha$ denotes a weighting parameter that balances minimizing the training loss and the validating loss for new data.
Equation (3) ensures that the model of each cluster is well-trained (i.e., minimizing the loss function during training) and that new data is accurately assigned to the most suitable cluster (i.e., minimizing the loss function during validating).
Figure 2: Framework of proposed model
@@ 4.1 Model architecture
In this paper, we propose CCRetNet based on two crucial components: Clustering and CNN-RetNet. The architecture of the proposed model is demonstrated in Fig 2. In brief, the aim of the proposed model is to qualify the uncertainty, extract features, and improve forecasting accuracy during the foggy winter.
We create the Tsallis Entropy by Weighted Permutation Pattern (TEWPP) to qualify the uncertainty. Based on this, hierarchical clustering is used to reduce the complexity. The hierarchical clustering utilizes the novel calculation of distance between series and the median linkage criterion as referenced in [34]. Additionally, we use modified RetNet to extract the features and forecast. CNN extracts the features, and RetNet forecasts PV power generation. NSGA-II optimizes the hyperparameters to enhance the robustness.
@@ 4.2 Clustering method
The section includes TEWPP and hierarchical clustering. We first introduce TEWPP following the model architecture.
@@ 4.2.1 The tsallis entropy by weighted permutation pattern
The Weighted Permutation Pattern is a method for time series analysis. It better captures complex structures compared to permutation patterns.
Considering series of PV power generation $\boldsymbol{x}=\{\boldsymbol{x}(1),$ $\boldsymbol{x}(2),\ldots,\boldsymbol{x}(T)\}$ with length $T$, we obtain its segment $\boldsymbol{x}^{m,\tau}_{t}=\{\boldsymbol{x}(t),\boldsymbol{x}(t+\tau),\ldots,\boldsymbol{x}(t+(m-1)\tau)\}$ for $t=1,2,\cdots$, where $m$ and $\tau$ respectively represent the dimension and time lag. Every $\boldsymbol{x}^{m,\tau}_{t}$ has $m!$ possible permutations, and it results in a unique permutation $\pi^{m,\tau}_{j}$.
Before giving the definition of the probability of weighted permutation pattern $p_{\omega}(\pi^{m,\tau}_{j})$, we give the definitions of $\bar{x}^{m,\tau}_{t}$ and weight value $\omega_{t}$.
@@
We create TEWPP applying the weighted permutation pattern to Tsallis entropy inspired by [36]. The calculation of TEWPP is defined as follows:
@@
We categorize the series of PV power into the cluster of large uncertainty and the cluster of small uncertainty based on TEWPP. Hierarchical clustering is used for the cluster of large uncertainty and the cluster of small uncertainty, respectively.
@@ 4.2.2 Hierarchical clustering
The core of hierarchical clustering is to construct a structure based on the distance between series and the distance between clusters. Therefore, different measures of distant calculation can lead to different clustering results. Next, we give the calculations of the distance between series (equation (10)) and the distance between clusters (equation (11)), the process of hierarchical clustering (algorithm 1) and the evaluation metric (equation (14)).
$\boldsymbol{x}_{p}=\{\boldsymbol{x}_{p}(t)\}^{T}$ and $\boldsymbol{x}_{q}=\{\boldsymbol{x}_{q}(t)\}^{T}$ are series of PV power generation with T-dimension for $p\neq q$. The weight $\boldsymbol{w}=\{w_{n}\}^{T}$ between the $\boldsymbol{x}_{p}$ and $\boldsymbol{x}_{q}$ is defined as below:
@@
We define the calculation for the distance between series. The calculation adopts the average of the $\{d^{\prime}_{n}\}^{T}$ in the middle, eliminating the influence of extreme values. Hence, the distance between $\boldsymbol{x}_{p}$ and $\boldsymbol{x}_{q}$ is defined as follows:
@@
The traditional distant calculations between different clusters for hierarchical clustering are single linkage criterion, complete linkage criterion, average linkage criterion, and median linkage criterion. We use a novel linkage criterion as referenced in [34] to compute the distance between different clusters as follows:
@@
where $C_{p}$ and $C_{q}$ stand for different clusters. $N=|C_{p}||C_{q}|$, where $|C_{p}|$ means the number of elements in cluster $C_{p}$. $r_{(1)}\leq r_{(2)}\leq\ldots\leq r_{(N)}$ denotes a rank-ordered set of the distances $d(\boldsymbol{x}_{p},\boldsymbol{x}_{q})$. Obviously, $r_{(1)}$ and $r_{(N)}$ represent the minimum distance and the maximum distance, respectively. To more effectively group similar segments while minimizing the impact of extreme values of the segments, the parameter $\alpha_{1}(N)$ and $\alpha_{2}(N)$ are set $N/2$ and $43N/$ 4 (i.e. $r_{\alpha_{1}(N)}$ and $r_{\alpha_{2}(N)}$ represent the second quartile $Q_{2}$ and the third quartile $Q_{3}$ of the distances $d(\boldsymbol{x}_{p},\boldsymbol{x}_{q})$ respectively).
We use hierarchical clustering according to the definitions of the above two distant calculations. The process of hierarchical clustering is shown in algorithm 1.
@@
We employ the silhouette coefficient [37] to determine the most suitable number of clusters. The silhouette coefficient is a way to evaluate the clustering effect, which combines cohesion and separation. The value spans from $-1$ to $1$. The closer silhouette coefficient of $\boldsymbol{x}_{p}$ is to 1, the more compact $\boldsymbol{x}_{p}$ is within its cluster.
Let $a(\boldsymbol{x}_{p})$ denotes the cohesion of the cluster $C_{p}$ and $b(\boldsymbol{x}_{p})$ denotes the separation between clusters. The calculations are defined as follows:

### v2

@@ Abstract
Solar energy is one of the most promising renewable energy resources. Forecasting photovoltaic power generation is an important way to increase photovoltaic penetration. However, the difficulty in qualifying uncertainty of PV power generation, especially during hazy weather, makes forecasting challenging. This paper proposes a novel model to address the issue. We introduce a modified entropy to qualify uncertainty during hazy weather while clustering and attention mechanisms are employed to reduce computational costs and enhance forecasting accuracy, respectively. Hyperparameters were adjusted using an optimization algorithm. Experiments on two datasets related to hazy weather demonstrate that our model significantly improves forecasting accuracy compared to existing models.
@@ 1 Introduction
Fossil fuels extracted for decades to meet global energy demands, causing serious environmental problems [1]. In contrast, solar power is known for its clean, pollution-free, and sustainable benefits [2]. Fig. 1 shows a significant increase in solar energy capacity over the past decade, with data derived from the International Renewable Energy Agency [3].
@@
Photovoltaic (PV) power generation embraces the property of uncertainty due to external factors [4]. Consequently, this uncertainty decreases forecasting accuracy. The power system may lead to serious failures if inaccurate PV forecasting is applied [5]. Consequently, accurate forecasting is essential to improve the stability of the power system [6].
@@
Existing PV forecasting models can be classified into physical, statistical, and hybrid models [7]. Physical models focus on theoretically explaining the works of PV power systems. Statistical models use data analysis to make forecasts. Hybrid models combine multiple models to improve accuracy. Table 1 lists existing PV forecasting models, showing their descriptions, advantages, and disadvantages.
[…]
@@
Existing works consider the impact of haze on PV forecasting [11]. Entropy is a way of qualifying uncertainty in time series [12]. We create modified entropy to qualify uncertainty affected by haze. Then, we use the clustering method based on modified entropy to reduce computational costs. Lastly, we use attention mechanism to make forecasts. Drawing from the analysis above, the main contributions of this work are presented as follows:
@@
  • We create the Tsallis Entropy by Weighted Permutation Pattern (TEWPP) to qualify uncertainty during hazy weather. TEWPP examines the distribution of patterns to qualify uncertainty. This method reflects the distribution of PV power generation during hazy weather.
@@
  • We use hierarchical clustering based on entropy to reduce the computational costs of the model. This strategy can improve the similarity of elements in the same cluster, thereby reducing the computational costs.
@@
  • We use modified Retention Network (RetNet) to extract the features and make forecasts. CNN uses kernel on input data to extract local features through multiplication and summation. RetNet is used to make forecasts. The model improves the forecasting accuracy.
@@
  • We adopt the Non-dominated Sorting Genetic Algorithm-II (NSGA-II) to optimize the hyperparameters. NSGA-II searches for optimal solutions by simulating natural selection and genetic mechanisms. The robustness of the model can be enhanced by using NSGA-II.
@@
  • We use datasets from a PV power station in a specific region of Jiangsu, China and another from Beijing, China. The experimental results are evaluated through evaluation metrics, and these results are compared to others obtained from traditional models.
@@
The remainder of this paper is organized as follows. Section 2 presents an overview of related works. Section 3 introduces the optimal goal of the system model. Section 4 gives an architecture and a detailed description of CCRetNet. Section 5 provides an overview of the experimental setup, air quality analysis, and the results of single-step and multi-step forecasting experiments in Jiangsu Province and Beijing. Section 6 provides conclusions.
@@
Physical models Simulate the physical process of PV power plants to make forecasts. Provide high-accuracy forecasts. Require large computational costs for measuring and updating data.
@@
Hybrid models Combine multiple models according to the characteristics of forecasting models to improve accuracy and reliability. Overcome limitations of a single model, improving accuracy and reliability. Result in higher computational costs than a single model.
@@ 2 Related works
Weather factors, particularly haze, are taken into account in the field of PV forecasting [17]. Yao et al. [18] introduced the air quality index and used the support vector machine algorithm to develop a new model for estimating solar radiation under haze conditions. The model improved accuracy and was suitable for regions with limited equipment or high pollution but did not explore using other neural network algorithms to assess the coupling effects of haze and weather parameters. Fan et al. [19] employed a support vector machine model to improve global and diffuse solar radiation predictions by incorporating air quality index and various pollutants. The approach significantly enhanced accuracy, particularly in heavily polluted areas. However, as the number of input pollutants increases, the complexity and computational resource consumption also rise, with diminishing returns on improvement. The research developed a new daily diffuse solar radiation model by incorporating the air quality index to improve accuracy in smog-affected areas [20]. The air quality index adjustment significantly enhanced accuracy, especially in polluted regions like Beijing. However, the model’s effectiveness varied across different regions, requiring further optimization for broader applicability.
Nowadays, the attention mechanism is used in the field of PV forecasting [21]. The research used improved complementary empirical mode decomposition and variational mode decomposition for reducing noise [2]. The attention mechanism focused on critical information to improve accuracy. The model was adjusted to data features to improve accuracy, whereas the complexity of the model limited its application. Yu et al. [13] used temporal, frequency, and fourier attention to capture detailed time-series insights from PV power data. The attention mechanism focused on crucial temporal and environmental features, thereby enhancing accuracy. However, this approach also obscured the significance of the time series itself. Wang et al. [22] combined a temporal convolutional network with an improved deep residual shrinkage network. The special attention mechanism in the model enhanced the ability to extract features in noisy environments. However, the model required significant computational resources.
Table 2 summarizes other related works, including details of works, advantages, and disadvantages, where "Ref." denotes "References".
@@
PV forecasting considering weather factors
@@
[11] 2021 Photovoltaic power; Power forecasting; Hazy conditions; Artificial neural network Enhanced accuracy in polluted environments. Relied on accuracy of air quality index data.
[23] 2020 Air pollution; Prediction; Solar radiation Improved accuracy with the bat algorithm, enhancing convergence speed. Resulted in larger computation under different pollution levels.
@@
[24] 2023 Photovoltaic power forecast; Channel attention mechanism; Time series forecasting Improved accuracy. Available for multivariate time series forecasting. Caused a risk of overfitting.
[25] 2023 Solar photovoltaic power; Short-term prediction; Deep learning; Attention net Enhanced processing capabilities through attention mechanisms. Sensitive to quality of input data.
[26] 2023 Temporal convolutional attention neural networks; PV power forecasting Captured the spatio-temporal features and improved the accuracy. Only represented the ultra-short-term prediction.
@@
MSR Multi-Scale Retention PV Photovoltaic
CCRetNet Clustering and CNN-RetNet Ref. Reference
CNN Convolutional Neural Network RetNet Retention Network
FLOPs Floating Point Operations RMSE Root Mean Squared Error
FFN Feed-Forward Network RNN Recurrent Neural Network
GRU Gate Recurrent Unit RS Random Search
LSTM Long Short Term Memory TPE Tree-structured Parzen Estimator
MAE Mean Absolute Error TEWPP The Tsallis Entropy by Weighted Permutation Pattern
@@ 3 Problem definition
Let $k$ represent the optimal number of clusters for PV power generation. $\mathcal{L}_{n}(\cdot)$ denotes the loss function of the forecasting model for the $n$-th cluster. The set of parameters is denoted by $\phi_{n}$, and the set of hyperparameters is denoted by $\psi_{n}$ for the $n$-th model. $\Phi=\{\phi_{1},\phi_{2},\cdots,\phi_{k}\}$ denotes the set of parameters for all model and $\Psi=\{\psi_{1},\psi_{2},\cdots,\psi_{k}\}$ denotes the set of hyperparameters for all model.
During the training phase, the objective is to minimize the loss function for each cluster. The objective function denotes as follows:
@@
The optimal set $\phi_{n}^{*}$ is determined during the training phase when equation (1) obtains its minimum value (i.e.,$\phi_{n}^{*}=\arg\min_{\Phi}\sum_{n=1}^{k}\mathcal{L}_{n}(\phi_{n})$).
For the validating phase, it should be assigned to the cluster that minimizes the loss function for the new data segment denoted by $\boldsymbol{x}$. The objective function denotes as follows:
@@
The optimal set $\psi_{n}^{*}$ is determined during the validating phase when equation (2) obtains its minimum value (i.e., $\psi_{n}^{*}=\arg\min_{\Psi}\sum_{n=1}^{k}\mathcal{L}_{n}(\psi_{n})$).
@@
where $\alpha$ denotes a weighting parameter that balances minimizing the training loss and the validating loss.
Equation (3) ensures that the model of each cluster is well-trained (i.e., minimizing the loss function during training) and that new data is assigned to the most suitable cluster (i.e., minimizing the loss function during validating).
@@ 4.1 Model architecture
In this paper, we propose CCRetNet based on two crucial components: Clustering and CNN-RetNet. The architecture of the proposed model is demonstrated in Fig. 2. In brief, the aim of the proposed model is to qualify uncertainty, extract features, and improve forecasting accuracy during hazy weather.
We create the Tsallis Entropy by Weighted Permutation Pattern (TEWPP) to qualify uncertainty. Based on TEWPP, hierarchical clustering is used to reduce the computational costs. The hierarchical clustering utilizes the novel calculation of distance between series and the median linkage criterion as referenced in [27]. Additionally, we use modified RetNet to extract the features and make forecasts. CNN extracts the features, and RetNet is used to make forecasts. NSGA-II optimizes the hyperparameters to enhance the robustness.
@@ 4.2 Clustering method
This section covers both TEWPP and hierarchical clustering. We first introduce TEWPP following the model architecture.
@@ 4.2.1 The tsallis entropy by weighted permutation pattern
The weighted permutation pattern is a method for time series analysis. Compared to permutation patterns, it better captures complex structures [28].
Considering a PV power generation series $\boldsymbol{x}=\{\boldsymbol{x}(1),$ $\boldsymbol{x}(2),\ldots,\boldsymbol{x}(T)\}$ with length $T$, we obtain its segment $\boldsymbol{x}^{m,\tau}_{t}=\{\boldsymbol{x}(t),\boldsymbol{x}(t+\tau),\ldots,\boldsymbol{x}(t+(m-1)\tau)\}$ for $t=1,2,\cdots$, where $m$ and $\tau$ respectively represent the dimension and time lag. Every $\boldsymbol{x}^{m,\tau}_{t}$ has $m!$ possible permutations, and it results in a unique permutation $\pi^{m,\tau}_{j}$.
Before defining the probability of the weighted permutation pattern $p_{\omega}(\pi^{m,\tau}{j})$, we first define $\bar{x}^{m,\tau}{t}$ and the weight value $\omega_{t}$.
@@
We create TEWPP by applying the weighted permutation pattern to Tsallis entropy inspired by [30]. The calculation of TEWPP is defined as follows:
@@
Based on TEWPP, we categorize the PV power series into clusters of large uncertainty and small uncertainty. Hierarchical clustering is used for the clusters of large uncertainty and small uncertainty, respectively.
@@ 4.2.2 Hierarchical clustering
The core of hierarchical clustering is to construct a structure based on the distance between series and the distance between clusters. Consequently, different distance measures can result in different clustering outcomes. We will next present the calculations for the distance between series (equation (10)) and between clusters (equation (11)), followed by the hierarchical clustering process (algorithm 1) and the evaluation metric (equation (14)).
Let $\boldsymbol{x}_{p}=\{\boldsymbol{x}_{p}(t)\}^{T}$ and $\boldsymbol{x}_{q}=\{\boldsymbol{x}_{q}(t)\}^{T}$ are series of PV power generation with T-dimension for $p\neq q$. The weight $\boldsymbol{w}=\{w_{n}\}^{T}$ between the $\boldsymbol{x}_{p}$ and $\boldsymbol{x}_{q}$ is defined as below:
@@
We define the calculation for the distance between series. The calculation adopts the average of the $\{d^{\prime}_{n}\}^{T}$ in the middle, thereby reducing the influence of extreme values. Hence, the distance between $\boldsymbol{x}_{p}$ and $\boldsymbol{x}_{q}$ is defined as follows:
@@
Traditional distance calculations between clusters in hierarchical clustering include the single linkage criterion, complete linkage criterion, average linkage criterion, and median linkage criterion. We use a novel linkage criterion as referenced in [27] to compute the distance between different clusters as follows:
@@
where $C_{p}$ and $C_{q}$ stand for different clusters. $N=|C_{p}||C_{q}|$, where $|C_{p}|$ means the number of elements in cluster $C_{p}$. $r_{(1)}\leq r_{(2)}\leq\ldots\leq r_{(N)}$ denotes a rank-ordered set of the distances $d(\boldsymbol{x}_{p},\boldsymbol{x}_{q})$. Obviously, $r_{(1)}$ and $r_{(N)}$ represent the minimum distance and the maximum distance, respectively. To more effectively group similar segments while minimizing the impact of extreme values of the segments, the parameter $\alpha_{1}(N)$ and $\alpha_{2}(N)$ are set to $N/2$ and $43N/$ 4 (i.e. $r_{\alpha_{1}(N)}$ and $r_{\alpha_{2}(N)}$ represent the second quartile $Q_{2}$ and the third quartile $Q_{3}$ of the distances $d(\boldsymbol{x}_{p},\boldsymbol{x}_{q})$ respectively).
We use hierarchical clustering according to the definitions of the above two distant calculations. The process of hierarchical clustering is detailed in algorithm 1.
@@
We employ the silhouette coefficient [31] to determine the most suitable number of clusters. The silhouette coefficient is a way to evaluate the clustering effect by combining cohesion and separation. The value spans from $-1$ to $1$. The closer silhouette coefficient of $\boldsymbol{x}_{p}$ is to 1, the more compact $\boldsymbol{x}_{p}$ is within its cluster.
Let $a(\boldsymbol{x}_{p})$ denotes the cohesion of the cluster $C_{p}$ and $b(\boldsymbol{x}_{p})$ denotes the separation between clusters. $a(\boldsymbol{x}_{p})$ is defined as mean distance from the segment $\boldsymbol{x}_{p}$ to all other segments in the same cluster. $b(\boldsymbol{x}_{p})$ denotes smallest mean distance from the segment $\boldsymbol{x}_{p}$ all segments in any other cluster. The calculations of $a(\boldsymbol{x}_{p})$ and $b(\boldsymbol{x}_{p})$ are defined as follows: